Q: A consumer's utility function is given by the expression: = (0.6X0.5+ 0.4Y0.5)2.

  • Determine the marginal utility functions for each commodity. Does marginal utility decrease when consumption increases?
  • Assuming that the price of good is Rs 15 and the price of Y is Rs 6, write the equation of the budget line and plot it when income is Rs 450. What is its slope? What does it indicate?
  • Calculate the marginal rate of substitution of Y for X and interpret its economic meaning. Write the equation showing the consumer's equilibrium condition.
  • Obtain the equilibrium values of X and Y.
  • Find the expressions for change in MUx due to an increase in Y and change in MUy due to an increase in X.
1
Expert's answer
2021-03-05T11:11:38-0500

a) U="(0.6X"0.5+0.4Y0.5)

MUx= "\delta"U/"\delta"x="2(0.6X"0.5+0.4Y0.5)"*0.3X"-0.5 =0.6X-0.5(0.6X0.5+0.4Y0.5)

MUy= "\delta"U/"\delta"y="2(06X"0.5+0.4Y0.5)*0.2Y-0.5 = 0.4Y-0.5 (0.6X0.5+0.4Y0.5)

The marginal utility decreases as the consumption increases.

b) Budget Line

PxX+PyY=M

15X+6Y=450

Slope =Py/Px = "6\/15 =0.4"

An increase in consumption of good Y leads to a decrease of good X by 0.4 units.

c) Marginal rate of Substitution MRS=("\delta"U/"\delta"y)/("\delta"U/"\delta"X)

=0.4Y-0.5(0.6X0.5+0.4Y0.5)/0.6X0.5(0.6X0.5+0.4Y0.5)

=0.4Y-0.5/0.6X0.5

=0.67X0.5Y-0.5

It implies that one good is substituted at a rate 0.67X0.5Y-0.5 such that if one increases, the other one will decrease at the same rate.

Consumer's equilibrium conditions

Y* = (0.67X0.5"Py\/Px)"0.5 X*=(0.67Y0.5"Py\/Px)"0.5